OpenStudy (anonymous):

The interest on an investment varies directly as the rate of interest. If the interest is ​$43 when the interest rate is 5.%, find the interest when the rate is 9.3​%.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (anonymous):

@radar

OpenStudy (rushwr):

Do u have the answer to check if I'm right or wrong ?

OpenStudy (anonymous):

Is this another one of those k=x/y problems?

OpenStudy (anonymous):

No. It's not multiple choice.

OpenStudy (anonymous):

What did you get? I got 76 but I think I did something wrong.

OpenStudy (rushwr):

I got $79.98

OpenStudy (anonymous):

That was right! How did you get the answer?

OpenStudy (rushwr):

Really?

OpenStudy (anonymous):

Yeah really

OpenStudy (rushwr):

oh okai If we take the investment as "x" Then , \[\frac{ 43 }{ x } * 100 = 5%\]

OpenStudy (rushwr):

right

OpenStudy (rushwr):

\[\frac{ interest }{ investment } * 100 = rate\]

OpenStudy (rushwr):

just tell me if u get it or not !

OpenStudy (anonymous):

I get it. And thanks!

OpenStudy (rushwr):

okai !!! So u get hw to solve right?